OpenStudy (anonymous):

It takes a runner 3 h, 1 min, 20 s to run a marathon. If the distance of a marathon is 42.2 km, what is the average speed of the runner?

OpenStudy (shane_b):

All you need to do is use the formula (converting the time as needed):\[Speed_{avg}=\frac{distance}{time}\]

OpenStudy (anonymous):

please atleast plug the number in the formula so i can get an understanding

OpenStudy (shane_b):

\[Speed_{avg}=\frac{42.2km}{3+\frac{1.333}{60}hrs}\]

OpenStudy (anonymous):

thanks

OpenStudy (shane_b):

np